OpenStudy (ikou):

"Solve 6 over x minus 4 equals 4 over x for x and determine if the solution is extraneous or not." I only need to find out if it is extraneous or not.

OpenStudy (ikou):

x = −8, extraneous x = −8, non-extraneous

OpenStudy (ikou):

I just need to know which one's correct

OpenStudy (ikou):

Please someone help me, I refuse to go to sleep until I finish this last question and its 3:30 AM

OpenStudy (ikou):

Everyone passes through this question thread.. It's four. This is ridiculous.

OpenStudy (ikou):

Please, I NEED to get some damn sleep. Somebody just help me..

OpenStudy (cderaalien):

lmao "some damn sleep." I feel you.

OpenStudy (cderaalien):

Lemme see what I can do.

OpenStudy (ikou):

Please help me kind sir

OpenStudy (ikou):

I will fan and medal

OpenStudy (cderaalien):

lol I am a chick but sure hold on

OpenStudy (ikou):

oh, please help me fair madam

OpenStudy (cderaalien):

I have your answer

OpenStudy (cderaalien):

I gave a link just in case you needed to explain the answer

OpenStudy (ikou):

but..

OpenStudy (ikou):

I've already been to that link

OpenStudy (ikou):

I'm dying inside

OpenStudy (cderaalien):

oh bloody hell,.... what type of class is this? I am only in Pre-Algebra.

OpenStudy (ikou):

Algebra II

OpenStudy (cderaalien):

Well, I looked at the link and it says the answer is negative 1/2

OpenStudy (cderaalien):

it's pretty simple to read

OpenStudy (ikou):

butbut, I only need to find out if its extraneous or not

OpenStudy (cderaalien):

Even though, I am not inside Algebra ll...something tells me

OpenStudy (cderaalien):

it is..extraneous

OpenStudy (ikou):

I am trusting you fully

OpenStudy (cderaalien):

I looked up the definition so.

OpenStudy (cderaalien):

Are you in FLVS?

OpenStudy (ikou):

It was wrong

OpenStudy (ikou):

and yess

OpenStudy (cderaalien):

pellet, I am sorry....I haven't even completed Algebra l yet. I'm sorry. Ugh, flvs...the devil. Good luck!

OpenStudy (ikou):

Es oki, too tired to care. It's totally fine dont worry, you did all you could.

OpenStudy (cderaalien):

alright anyway bye :/

OpenStudy (anonymous):

x = −8, non-extraneous http://www.mathsisfun.com/definitions/extraneous-root.html

OpenStudy (cderaalien):

@robtobey thank you!!

OpenStudy (anonymous):

\[\frac{6}{x-4}=\frac{4}{x} \]

OpenStudy (ikou):

Im going to have a screaming fit..

OpenStudy (ikou):

Anywho, I am going to go to sleep for 5 hours..